# Hyper-Connections

### v1

@@ Acknowledgements
This research was conducted at ByteDance Inc. We are grateful for the suggestions and assistance provided by Yaowei Zheng, Yuyu Zhang, Yunshui Li, Xiang Li, and Bairen Yi.
Appendix A MoE 1B/7B Model Experiments
[…]
Appendix B 7B Model Experiments
[…]
Appendix C Vision Experiments
@@ C.1 Image Generation
[…]
Table 7: Benchmarking class-conditional image generation on ImageNet $\times256$, with cfg=1.50. NP, P, and R are short for Numerical Precision, Precision, and Recall, respectively.
@@
C.2 Image Classification
For the image classification experiments, we train ViT/16-Base and ViT/16-Large models with images at a resolution of $224\times 224$ for 300 epochs, following the experimental setup used by (Dosovitskiy et al., 2020).To speed up the training process, we use bfloat16 numerical precision. The training configuration is detailed in Table 9. Within this configuration, we replace the residual connections with static and dynamic hyper-connections, referred to as SHC and DHC, respectively, using an expansion rate of $n=2$. The top-1 accuracy results are presented in Table 8.
@@ C.2 Image Classification
Table 9: Training hyperparameters for ViT.
Appendix D More Visualization and Analysis
@@
[…]
Appendix E Derivation of Non-Trainable Hyper-Connection Matrix for Residual Connections
E.1 Pre-Norm Residual Connection
@@
E.2 Post-Norm Residual Connection
@@
where $\mathcal{A}$ is the affine transformation, and $\mathcal{C}$ is the re-centering operator. Thus, the mean of the output of $\mathcal{T}$ is 0.
@@
Appendix F Sequential-Parallel Duality
F.1 Hyper-Connection Matrix of Sequential Arrangement
@@
F.2 Hyper-Connection Matrix of Parallel Arrangement
@@
Appendix G Pseudocode of Hyper-connections
@@
[…]
@@
8:  $\begin{pmatrix}\mathbf{h}_{0}&\mathbf{H^{\prime}}\end{pmatrix}\leftarrow\mathcal{WC}^{k\intercal}\mathbf{H}$ $\triangleright$ Width Connections
9:  $\mathbf{h}^{\prime}_{0}\leftarrow\mathcal{T}^{k}(\mathbf{h}_{0})$ $\triangleright$ Layer Computation
10:  $\mathbf{\hat{H}}\leftarrow\mathbf{B}^{k\intercal}\mathbf{h}^{\prime}_{0}+\mathbf{H^{\prime}}$ $\triangleright$ Depth Connections
@@
[…]
@@
Appendix H PyTorch Implementation of Hyper-connections
@@
class HyperConnection($n$.Module):
@@
self.static_beta = $n$.Parameter(torch.ones((rate,), device=device))
@@
self.static_alpha = $n$.Parameter(torch.cat([init_alpha0, torch.eye((rate), device=device)], dim=1))
@@
self.dynamic_alpha_fn = $n$.Parameter(torch.zeros((dim, rate+1), device=device))
self.dynamic_alpha_scale = $n$.Parameter(torch.ones(1, device=device) * 0.01)
self.dynamic_beta_fn = $n$.Parameter(torch.zeros((dim, ), device=device))
self.dynamic_beta_scale = $n$.Parameter(torch.ones(1, device=device) * 0.01)
@@
Appendix I Validation Sets and Downstream Tasks
@@
Table 11: Downstream Benchmarks for OLMoE.
Appendix J 1B Model Experiments
Method arc_easy copa hellaswag openbook_qa piqa sciq winogrande avg
@@
Table 12: Results on Downstream Benchmarks for 1B Model.
@@
Table 13: Losses of V2 validation sets for 1B Model.
@@
Table 14: Perplexities of V2 validation sets for 1B Model.
@@
Table 15: Losses of V3 validation sets for 1B Model.

### v2

@@ Acknowledgements
This research was conducted at ByteDance Inc. We are grateful for the suggestions and assistance provided by Yaowei Zheng, Yuyu Zhang, Yunshui Li, Xiang Li, Bairen Yi, Zhenyi Lu and Xintian Han.
Appendix A Transformer with Hyper-Connections
Figure 8: Comparison between transformers with hyper-connections and that with residual connections.
Appendix B Parameters, Computation and Memory Footprint Analysis
Static Hyper-Connections. All learnable parameters are included in the hyper-connection matrix $\mathcal{HC}$ in Eq. 1. The number of parameters in one $\mathcal{HC}$ is given by:
$\left|\theta_{\texttt{SHC}}\right|=|\theta_{\mathbf{B}}|+|\theta_{\mathbf{A}}|=n+n\cdot(n+1)=n\cdot(n+2),$ (21)
where $n$ is the expansion rate, $|\theta_{\mathbf{B}}|$ is the number of parameters in $\mathbf{B}$ in SHC, and $|\theta_{\mathbf{A}}|$ is the number of parameters in $\mathbf{A}$. Each layer contains two hyper-connection modules (one for the self attention and one for the feedforward network). Thus, the number of extra parameters is:
$P_{\texttt{extra}}=\left|\theta_{\texttt{SHC}}\right|\times 2\times L,$ (22)
where $L$ is the number of layers. For example, in OLMo-1B-SHC×$\times$4, $P_{\texttt{extra}}=4\times(4+2)\times 2\times 16=768$.
Dynamic Hyper-Connections. The parameters of DHC are defined in Eqs. 10, 11, 12, and 13, and the number of parameters is given by:
$\left|\theta_{\texttt{DHC}}\right|$ $=|\theta_{\texttt{norm}}|+|s_{\beta}|+|\theta_{\mathbf{W}_{\beta}}|+|\theta_{\mathbf{B}}|+|s_{\alpha}|+|\theta_{\mathbf{W}_{m}}|+|\theta_{\mathbf{A}_{m}}|+|\theta_{\mathbf{W}_{r}}|+|\theta_{\mathbf{A}_{r}}|$ (23)
$=|\theta_{\texttt{norm}}|+1+d_{\texttt{model}}+n+1+d_{\texttt{model}}+n+d_{\texttt{model}}\times n+n\times n$ (24)
$=|\theta_{\texttt{norm}}|+d_{\texttt{model}}\times(n+2)+n\times(n+2)+2,$ (25)
where $d_{\texttt{model}}$ is the dimension of the hidden states in the transformer, and $|\theta_{\texttt{norm}}|$ depends on the type of normalization module. In OLMo models, there are no parameters for normalization, so $|\theta_{\texttt{norm}}|=0$. In OLMoE, $|\theta_{\texttt{norm}}|=d_{\texttt{model}}$. Similar to the static hyper-connections, the number of extra parameters is:
$P_{\texttt{extra}}=\left|\theta_{\texttt{DHC}}\right|\times 2\times L,$ (26)
For example, for OLMo-1B-DHC×$\times$4, $P_{\texttt{extra}}=(0+2048\times(4+2)+4\times(4+2)+2)\times 2\times 16=394,048$.
The number of parameters for DHC and SHC used in the experiments is detailed in Table 7, while their corresponding FLOPs comparisons are provided in Table 8. Regardless of whether SHC or DHC is used, the additional parameters and computational overhead introduced are minimal and can be considered negligible.
Table 7: Comparison of number of parameters.
Method HC Params(B) Total Params(B) Total Params $\Delta$ rate (%)
OLMo-1B - 1.17676442 -
OLMo-1B-SHC×$\times$2 0.0000026 1.17676467 +0.00002%
OLMo-1B-SHC×$\times$4 0.0000077 1.17676518 +0.00007%
OLMo-1B-DHC×$\times$2 0.0002625 1.17702688 +0.02230%
OLMo-1B-DHC×$\times$4 0.0003940 1.17715846 +0.03349%
OLMo-7B - 6.88809574 -
OLMo-7B-DHC×$\times$4 0.0013124 6.88967027 +0.02286%
OLMoE-1B-7B - 6.91909427 -
OLMoE-1B-7B-DHC×$\times$4 0.0003940 6.91948832 +0.00570%
Computation Analysis. The main computational cost of SHC and DHC lies in line 5 of Algorithm 1, where the complexity is $\mathcal{O}(d_{\text{model}}\times n\times(n+1))$. The computational cost of the FFN is $\mathcal{O}(2\times d_{\text{model}}\times d_{\text{ffn}})$, and that of the projection part of attention is $\mathcal{O}(4\times d_{\text{model}}\times d_{\text{model}})$. Since $\mathcal{O}(d_{\text{model}}\times n\times(n+1))\ll\mathcal{O}(4\times d_{\text{model}}\times d_{\text{model}})<\mathcal{O}(2\times d_{\text{model}}\times d_{\text{ffn}})$, the computational cost of HC is negligible compared to the cost of both FFN and the attention projection part. Here, $d_{\text{ffn}}$ is the inner dimension of the FFN. The detailed computation cost statistics are presented in Table 8.
Table 8: FLOPs per token in forward pass.
Method HC FLOPs (G) Total FLOPs (G) Total FLOPs $\Delta$ rate (%)
OLMo-1B - 2.3536 -
OLMo-1B-SHC×$\times$2 0.0010 2.3545 +0.038%
OLMo-1B-SHC×$\times$4 0.0031 2.3566 +0.127%
OLMo-1B-DHC×$\times$2 0.0020 2.3554 +0.076%
OLMo-1B-DHC×$\times$4 0.0049 2.3583 +0.200%
OLMo-7B - 13.3647 -
OLMo-7B-DHC×$\times$4 0.0197 13.3844 +0.147%
OLMoE-1B-7B - 2.3580 -
OLMoE-1B-7B-DHC×$\times$4 0.0049 2.3629 +0.208%
Memory Footprint. The introduction of HC results in a minor increase in activation memory usage during training. For a transformer model with $L$ layers, a model dimension of $d_{\text{model}}$, batch size $b$, sequence length $s$, and number of attention heads $a$, the activation memory is calculated as $sbd_{\text{model}}L(34+5as/d_{\text{model}})$, as outlined in Korthikanti et al. (2022). Incorporating HC with an expansion rate of $n$ adds an extra memory overhead of $2nsbd_{\text{model}}L$. For $n=2$, this contributes less than $15\%$ to the total memory usage of a standard transformer. Notably, the memory consumption is mostly driven by the weight parameters, which experience only a slight increase with HC. Additionally, given HC’s low computational cost, the hidden states generated by HC can be discarded post forward pass and recomputed during backpropagation to further optimize memory usage. With this approach, the additional memory requirement is reduced to $nsbd_{\text{model}}$. During inference, the memory usage for activations is largely determined by the Key-Value cache, which is not impacted by the extra activations brought by HC. Moreover, the hidden states from earlier layers can be released as soon as the next layer’s computations start, significantly lowering memory requirements.
Appendix C MoE 1B/7B Model Experiments
Figure 9: Loss curves in V3 validation sets and accuracy curves on downstream tasks for OLMoE-1B7B and OLMoE-1B7B-DHC×$\times{4}$ models.
Appendix D 7B Model Experiments
Figure 10: Loss curves in V3 validation set and accuracy curves on downstream tasks for OLMo-7B and OLMo-7B-DHC×$\times$4 models.
Appendix E Vision Experiments
@@ E.1 Image Generation
Table 9: Benchmarking class-conditional image generation on ImageNet 256$\times$256, with cfg=1.50. NP, P, and R are short for Numerical Precision, Precision, and Recall, respectively.
@@
DiT-XL/2-SHC×$\times$2 FP16 675 2.18 4.52 287.24 0.82 0.60
@@
E.2 Image Classification
For the image classification experiments, we train ViT/16-Base and ViT/16-Large models with images at a resolution of $224\times 224$ for 300 epochs, following the experimental setup used by (Dosovitskiy et al., 2020).To speed up the training process, we use bfloat16 numerical precision. The training configuration is detailed in Table 11. Within this configuration, we replace the residual connections with static and dynamic hyper-connections, referred to as SHC and DHC, respectively, using an expansion rate of $n=2$. The top-1 accuracy results are presented in Table 10, and the training loss curves for ViT/16-Large and ViT/16-Large with DHC×$\times$2 are shown in Fig. 11.
@@ E.2 Image Classification
Figure 11: Training loss curves of ViT/16-Large and ViT/16-Large-DHC×$\times$2, smoothed using an Exponential Moving Average (EMA) with a decay rate of 0.999. The gain from Hyper-Connections decreases as training progresses, likely due to pass over the same dataset across many epochs, resulting in diminishing returns from the additional capacity provided by Hyper-Connections.
E.3 Visulization of DHC
We randomly select three categories from the ImageNet dataset and sample the corresponding examples from the validation set. These samples are fed into the ViT-Base/16-DHC×$\times$2 model to compute the dynamic connection weights of the DHC in the final layer. As shown in Fig. 12, we visualize the distribution of these weights. We observe that the intra-class distribution of beta is highly concentrated, indicating that samples within the same category tend to have similar beta values. In contrast, the distribution of alpha is less concentrated, but the differences between the distributions of different categories are more pronounced, as exemplified by $\alpha_{2,0}$.
Figure 12: Distribution of weights of last DHC in ViT-Base/16-DHC×$\times$2 model.
Table 11: Training hyperparameters for ViT.
@@ E.3 Visulization of DHC
Appendix F More Visualization and Analysis
@@
[…]
(a) OLMo-$\times 1$
(b) OLMo-1B-DHC×$\times 2$
(c) OLMo-1B-DHC×$\times 4$
Figure 14: Comparison of unfolded connection matrices for OLMo-$\times 1$, OLMo-1B-DHC×$\times 2$ and OLMo-1B-DHC×$\times 4$ model.
@@
We show the connection matrices for OLMo-1B-SHC×$\times$4 model in Fig. 13(b). Comparing to DHC, as shown in Fig. 13(a), SHC shares exactly the same connection patterns. Moreover, we observe many more PTB-like blocks in SHC, e.g., layers from 13 to 18. Note that the connection relation for SHC is token independent, and such PTB-like blocks can be physically reorganized to be parallelly computed.
How HC×$\times 1$ fails.
The OLMo-$\times 1$ model is observed to perform worse than baseline in our experiments. Its connection matrix is visualized in Fig. 14 to show how it fails. Above all, we observe that layer 17 is wasted, who has no connection to subsequent layers at all. Secondly, compared to HC×$\times 2$ and HC×$\times 4$ models, the $\Lambda$ shaped pattern does not appear. Note that HC×$\times 1$ does not support the pattern of $\Lambda$ in its mathematical formulation, where the connections to previous layers must be weakened or strengthened simultaneously. Thus, the lack of connection from the early layers to the final layers may suffer from gradient vanishing, like post-norm style transformers, which leads to performance degeneration.
Appendix G Derivation of Non-Trainable Hyper-Connection Matrix for Residual Connections
G.1 Pre-Norm Residual Connection
@@
G.2 Post-Norm Residual Connection
@@
where $\mathcal{A}$ is the affine transformation, and $\mathcal{C}$ is the re-centering operator. Thus, the mean of the output of $\mathcal{T}$ is 00.
@@
Appendix H Sequential-Parallel Duality
H.1 Hyper-Connection Matrix of Sequential Arrangement
@@
H.2 Hyper-Connection Matrix of Parallel Arrangement
@@
Appendix I Pseudocode of Hyper-connections
@@
5:$\mathbf{H}^{0}\leftarrow\begin{pmatrix}\mathbf{h}^{0}&\mathbf{h}^{0}&\dots&\mathbf{h}^{0}\end{pmatrix}^{\intercal}\in\mathbb{R}^{n\times d}$
@@
8:     $\begin{pmatrix}\mathbf{h}_{0}&\mathbf{H^{\prime}}\end{pmatrix}\leftarrow\mathcal{WC}^{k\intercal}\mathbf{H}$triangleright Width Connections
9:     $\mathbf{h}^{\prime}_{0}\leftarrow\mathcal{T}^{k}(\mathbf{h}_{0})$triangleright Layer Computation
10:     $\mathbf{\hat{H}}\leftarrow\mathbf{B}^{k\intercal}\mathbf{h}^{\prime}_{0}+\mathbf{H^{\prime}}$triangleright Depth Connections
@@
14:$\mathbf{h}^{L}\leftarrow\text{sum rows of }\mathbf{H}^{L}$
15:$\mathbf{h}^{L}\leftarrow\text{Normalization Layer}(\mathbf{h}^{L})$
@@
Appendix J PyTorch Implementation of Hyper-connections
@@
class HyperConnection(nn.Module):
@@
self.static_beta = nn.Parameter(torch.ones((rate,), device=device))
@@
self.static_alpha = nn.Parameter(torch.cat([init_alpha0, torch.eye((rate), device=device)], dim=1))
@@
self.dynamic_alpha_fn = nn.Parameter(torch.zeros((dim, rate+1), device=device))
self.dynamic_alpha_scale = nn.Parameter(torch.ones(1, device=device) * 0.01)
self.dynamic_beta_fn = nn.Parameter(torch.zeros((dim, ), device=device))
self.dynamic_beta_scale = nn.Parameter(torch.ones(1, device=device) * 0.01)
@@
Algorithm 3 Pseudocode of transformer with hyper-connections in a PyTorch-like style.
# h: hyper hidden matrix (BxLxNxD)
# atten_hyper_connection, ffn_hyper_connection: hyper-connection modules
# attn_norm, ffn_norm: normalization modules
# Attention Block
mix_h, beta = atten_hyper_connection.width_connection(h)
h = attn_norm(mix_h[...,0,:])
h = self_attention(h)
h = atten_hyper_connection.depth_connection(mix_h, dropout(h), beta)
# FFN Block
mix_h, beta = ffn_hyper_connection.width_connection(h)
h = ffn_norm(mix_h[...,0,:])
h = ffn(h)
h = ffn_hyper_connection.depth_connection(mix_h, dropout(h), beta)
Appendix K Validation Sets and Downstream Tasks
@@
Appendix L 1B Model Experiments
Figure 15: Training loss curves of related works, smoothed using Exponential Moving Average (EMA) with a decay rate of 0.99.
Figure 16: Training loss curves of DHC with tanh over 500 billion tokens, smoothed using Exponential Moving Average (EMA) with a decay rate of 0.99.
Figure 17: Training loss curves of DHC without tanh over 500 billion tokens, smoothed using Exponential Moving Average (EMA) with a decay rate of 0.99.
Table 14: Results on downstream benchmarks for 1B models.
Method arc_easy copa hellaswag openbook_qa piqa sciq winogrande avg.
@@
Table 15: Losses of V2 validation sets for 1B Model.
@@
Table 16: Perplexities of V2 validation sets for 1B models.
@@
Table 17: Losses of V3 validation sets for 1B model.
@@
Table 18: Perplexities of V3 validation sets for 1B models.